OpenStudy (anonymous):

Find the derivative of y=-xsinx when x=pie Find the derivative of y=e^(-2x) What is the coeffient of your answer. Find (d^2y)/(dx^2) when x=1 y=(x^3/3)+(x^2/2)+(x) Please explain, I don't really understand deivatives.

zepdrix (zepdrix):

So how bout that first one?\[\large\rm y=-x \sin x\]`Setting up` my product rule gives me\[\large\rm y'=\color{royalblue}{(-x)'}\sin x+(-x)\color{royalblue}{(\sin x)'}\]

zepdrix (zepdrix):

Take derivative of the blue parts. What do you get? :d (-x)' = ? (sin x)' = ?